OpenStudy (anonymous):

Candice needs a plumber and must decide between 2 companies. For a service visit, Company A charges $60 to send a plumber plus $45/h. Company B charges $70 to send a plumber plus $35/h. The graph that models this situation is shown here. According to the graph, what is the cost when the two plumbers have worked the same amount of time and cost is the same? $ Graphs of Cost vs Time for Company A and Company B. Cost is on the vertical axis and time is on the horizontal axis. The line for Company A goes through the points (0, 60), (1, 105), (1.5, 132.5). The line for Company B goes th

OpenStudy (anonymous):

C is cost. h is time in hours. A: C=60+45h B: C=70+35h You can graph both and find the intersection. The intersection is at h=1 hour.

OpenStudy (anonymous):

This means that the cost is the same at h=1 hour. Plug in 1 for h, and you yield $105 as the total cost.

MsBrains (ms-brains):

Hi! Welcome to OpenStudy! :)